# Investigating Feature Attribution for 5G Network Intrusion Detection

### v1

@@ 2.2 VoTE-XAI
Where $x_{1},\dots,x_{n}$ are the features, $c_{1},\dots,c_{n}$ are their respective values, and $d$ is the model’s prediction. An explanation $E$ is minimal if removing any element from $E$ invalidates the explanation. The formal definition is provided in Equation 3:
$\forall A\subset E,\bigwedge_{(x_{i},c_{i})\in A}(x_{i}=c_{i})\mathrel{{\ooalign{$\not\phantom{=}$\cr$\implies$}}}f(x_{1},\dots,x_{n})=d$ (3)
VoTE-XAI starts with a full explanation consisting of all feature-value pairs involved in a prediction. It then applies a deletion filter that iteratively removes features. An abstract interpretation-based oracle efficiently checks whether each reduced explanation still leads to the same prediction. If so, the feature is discarded. Otherwise, it is kept in the minimal set.
3 Investigating Feature Attribution for 5G Attacks
This study investigates the effectiveness of the above two representative methods in the context of 5G security monitoring. We conjecture that logic-based explanations, such as those provided by VoTE-XAI, offer a more concise, stable, and computationally efficient representation of feature contributions to classifier alerts (referred to as predictions above). This is particularly critical for real-time security applications, where delayed, overly verbose, or misleading explanations can hinder effective threat mitigation and timely response.
@@ 3.1 Research Questions
To test our initial conjecture, we define the following research questions:
@@
    RQ1: How do statistical (SHAP) and logic-based (VoTE-XAI) approaches differ in identifying and prioritizing critical features relevant for intrusion detection alerts across different 5G security scenarios?
@@
    RQ2: What are the computational efficiency trade-offs between SHAP and VoTE-XAI in generating explanations, and how do these trade-offs impact real-time 5G security monitoring?
@@ 3.2 Investigation Approach
The workflow is summarized in Figure 1. It consists of four steps which are further detailed in Section 5:
@@
    Raising alerts with Trained Model: An XGBoost classifier is trained and used to generate alerts on the test data. The model’s classification output serves as the input for subsequent explainability analysis.
@@
    Selection of True Positives for Explanation: Only true positive (TP) malicious flows are selected for explanation, ensuring the analysis focuses on correctly identified attacks.
@@
    Explanation Generation via Feature Attribution: For each alert, statistical explanations are derived using SHAP, and logic-based ones are computed using VoTE-XAI.
@@
    Analysis: The resulting explanations are evaluated based on the metrics defined in Section 5. The findings help evaluate the conciseness of the explanations, determine whether the two methods produce contradictory explanations, and assess computational efficiency.
@@ 4 Incident-response Case Studies
Open datasets for studying 5G network attacks and defenses are few and far between. The existing ones are not ideal due to a severe imbalance in data (benign/attack) categories since they are created in research labs for attack classification or the study of a specific mitigation response. However, our goal here is to create interpretations of alerts raised while monitoring, attributing the attack to identified feature sets. Therefore, potential deficiencies in the datasets will not be relevant for our work, which focuses on the alert level.
@@ 4.1 Dataset Description
We use three available datasets for studying 5G attacks: the 5G Network Intrusion Detection (5G-NIDD) [36], Multistep Attack (MSA) [37], and Packet Forwarding Control Protocol (PFCP) [38] datasets. Although we could have applied our approach to better-known but not 5G-related network intrusion detection benchmarks, we observe that those datasets are rather low-dimensional relative to the complexities in 5G networks reflected in terms of the number of features. For instance, CICIoT2023 [39] and CIC UNSW-NB15 [40] contain 46 and 76 features, respectively, which is fewer than even our use case with the lowest dimension in the next subsections. The details of the datasets adopted are in the following, and summarized in Table 1.
@@
The 5G-NIDD provides labeled network traffic data, distinguishing between normal operations and attack scenarios. It includes 92 network flow features, such as flow duration, packet size statistics, protocol types, flow flags, and packet inter-arrival times.

### v2

@@ 3.3 Logic-based Feature Attribution and VoTE-XAI
Where $x_{1},\dots,x_{n}$ are the features, $c_{1},\dots,c_{n}$ are their respective values, and $d$ is the model’s prediction [56]. An explanation $E$ is minimal if removing any element from a valid explanation $E$ invalidates it. The formal definition is provided in Equation 3:
$\forall A\subset E,\bigwedge_{(x_{i},c_{i})\in A}(x_{i}=c_{i})\mathrel{{\vtop{\halign{#\cr$\not\phantom{=}$\cr$\implies$\crcr}}}}f(x_{1},\dots,x_{n})=d$ (3)
Equations 2 and 3 have been defined in [56]. VoTE-XAI starts with a full explanation consisting of all feature-value pairs involved in a prediction. It then applies a deletion filter that iteratively removes features. An abstract interpretation-based oracle efficiently checks whether each reduced explanation still leads to the same prediction. If so, the feature is discarded. Otherwise, it is kept in the minimal set.
4 Method
The objective of this study is to get a deeper understanding of the effectiveness of the above two representative methods in the context of 5G security monitoring. We are interested in assessing whether the two methods perform similarly in terms of identified features in each attack scenario and across similar ones. In addition, we want to understand whether the timing performance of the two methods is appropriate in this 5G context. To be as fair as possible in the comparative analysis, we will use an efficient tree-specific version of SHAP in Section 4.2, Step 3.
@@ 4.1 Research Questions
To systematically compare the two methods, we define the following research questions:
@@
    RQ1: How do SHAP and VoTE-XAI differ in identifying and prioritizing critical features for a given alert in 5G security monitoring?
@@
    RQ2: How efficient are SHAP and VoTE-XAI in generating attributions, and to what extent does this impact real-time response and forensic analysis?
@@ 4.2 Investigation Approach
The method is summarized in Figure 1. It consists of four steps which are further detailed in Section 6:
@@
    Raising Alerts with Trained Model: An XGBoost classifier is trained and used to generate alerts on the test data. The model’s classification output serves as the input for subsequent steps.
@@
    Selection of Alerts for Explanation: The outputs of the model will be of two categories (negatives and positives). The True Negatives (TNs) are the huge majority of cases in a real scenario, and there is no reason nor way to explain them. They simply pass the monitoring quietly. False Negatives (FNs) correspond to missed threats, and the operator would not be able to act upon something that is not even brought to their attention. Among the positives, the True Positives (TPs) are the most urgent to act upon, and the main reason for attribution, hence the focus of our method. With respect to False Positives (FPs), attribution may be useful to decide to de-prioritize them [13]. The overarching goal of our work is to see how attribution may be useful for future mitigation actions. FPs cannot be mitigated, only identified and ignored. Hence, this Step will focus on TPs [^5] However, a comment about how the few FPs in our experiments would be treated when explained is included in section 6.1. Note that the number of such alerts is so small that a statistical evaluation would not be meaningful..
@@
    Explanation Generation via Feature Attribution: For each alert, statistical attributions are derived using SHAP TreeExplainer optimization, and logic-based ones are computed using VoTE-XAI. At this stage, the feature attribution engines take as input both the selected alerts and the XGBoost model. This is necessary since the methods need to access the internal structure of the model to have the context within which alerts are raised.
@@
    Analysis: The resulting attributions are evaluated based on the metrics defined in Section 6. The findings help evaluate the sparsity and stability of the attributions, determine the degree of alignment between the two methods, and assess computational efficiency.
@@ 5 Incident-response Case Studies
Open datasets for studying 5G network attacks and defenses are few and far between. The existing ones are not ideal due to a severe imbalance in data (benign/attack) categories since they are created in research labs for attack classification or the study of a specific mitigation response. They contain far more attacks than benign data compared to real world. However, our goal here is to create interpretations of alerts raised while monitoring, attributing the attack to identified feature sets. Therefore, potential deficiencies in the datasets will not be relevant for our work, which focuses on the alert level.
@@ 5.1 Dataset Description
We use three available datasets for studying 5G attacks: the 5G Network Intrusion Detection (5G-NIDD) [48], Multistep Attack (MSA) [38], and Packet Forwarding Control Protocol (PFCP) [2] datasets. Although we could have applied our approach to better-known but not 5G-related network intrusion detection benchmarks, we observe that those datasets are rather low-dimensional relative to the complexities in 5G networks reflected in terms of the number of features. For instance, CICIoT2023 [40] and CIC UNSW-NB15 [36] contain 46 and 76 features, respectively, which is fewer than even our use case with the lowest dimension in the next subsections.
All selected datasets have been preprocessed by removing duplicate entries and encoding labels using sklearn LabelEncoder, following standard ML pipeline procedures. Then, the flows were split into training and test subsets using an 80:20 ratio and preserving the class proportions of the full datasets.
The details of the adopted datasets are in the following, and summarized in Table 2.
Table 2: Summary of datasets’ dimensionality, class balance, and train/test splits.
Dataset # Features Total Samples Train Set Test Set # Attack Types
Benign Malicious Benign Malicious
5G-NIDD [48] 92 1,215,890 381,904 590,812 95,833 147,341 8
MSA [38] 478 25,160 719 19,409 186 4,846 21
PFCP [2] 83 7,199 1,028 4,114 411 1,646 4
@@
The 5G-NIDD provides labeled network traffic data, distinguishing between normal operations and attack scenarios. It includes 92 network flow features, such as flow duration, packet size statistics, protocol types, flow flags, and packet inter-arrival times. The original documentation was not complete in terms of explaining the features of the data set. Thus, additional work was needed to understand the meaning of the feature names referred to. These are useful for understanding the rest of this paper. Table 3 presents a subset of the most relevant network flow features identified across the methods applied in this study.